### v1

# Conditional Dependence via U-Statistics Pruning

Abstract
The problem of measuring conditional dependence between two random phenomena arises when a third one (a confounder) has a potential influence on the amount of information shared by the original pair. A typical issue in this challenging problem is the inversion of ill-conditioned autocorrelation matrices. This paper presents a novel measure of conditional dependence based on the use of incomplete unbiased statistics of degree two, which allows to re-interpret independence as uncorrelatedness on a finite-dimensional feature space. This formulation enables to prune data according to the observations of the confounder itself, thus avoiding matrix inversions altogether. Moreover, the proposed approach is articulated as an extension of the Hilbert-Schmidt independence criterion, which becomes expressible through kernels that operate on 4-tuples of data.
Index Terms:
Hilbert-Schmidt independence criterion (HSIC), kernel methods, co-information, U-Statistics, conditional dependence.
I Introduction
The problem of measuring statistical dependence emerges in many fields, such as multivariate statistical analysis [1, Ch. 6], wireless sensor networks [2, Ch. 2], data analytics [3], and statistical array processing [4]. The Hilbert-Schmidt independence criterion (HSIC[5] provides a versatile non-negative measure of statistical dependence that becomes null if and only if the data pair is independent. Grounded in the theory of reproducing kernel Hilbert space s (RKHS), the HSIC succeeds in discovering both linear and nonlinear relationships in data. A well-known interpretation of the HSIC is that dependence produces a correlation of distances [6], i.e. close couples in one data set coincide with close couples in the other data set. Another insightful interpretation is that the HSIC measures statistical dependence by gauging the correlation of data implicitly mapped onto an infinite-dimensional space. This basic idea of translating dependence to correlation also admits a more traditional implementation by explicitly mapping data onto a finite-dimensional space [7], with the advantage of leveraging classical second-order statistics for inferring information-theoretic measures.
Conditional dependence becomes relevant when a third phenomenon might explain, mediate, or confound the apparent relationship exhibited by the original random pair, which happens in fields such as causal discovery and Bayesian network learning [8]. Although mapping data onto a space of higher dimension is still useful for that purpose, it is still necessary to measure conditional correlation in this newfound space. This method usually leads to the computation of the Schur complement of high-dimensional matrices [9, Ch. 4], which is known to have numerical issues concerned with matrix inversions and regularization due to the usually low rank structure of the involved autocorrelation matrices [10].
This letter presents a procedure for measuring conditional dependence by statistically conditioning the observed data to a potential confounder, which bypasses the matrix inversion problems. First, a novel theoretical derivation of the classical HSIC is provided in Sec. II, by translating the problem of statistical dependence into one of correlation after mapping the data onto finite-dimensional spaces based on steering vectors [7]. Sec. III briefly reviews the theory of unbiased statistics (U-Statistics) [11], which is then employed in Sec. IV to show that conditional dependence can be accomplished by just pruning U-Statistics under the control of the confounder [12]. The obtained measure, the conditional HSIC (C-HSIC), is based on a signal processing structure resembling kernel methods, which embraces the HSIC as a particular case when the U-Statistic is complete. However, it has the distinctive feature of being articulated on 4-tuples instead of the classical processing based on pairs of data from each phenomenon.
II Marginal dependence as correlation
Let $\mathbf{d}:\mathbb{R}\to\mathbb{C}^{M}$ be a mapping based on windowed steering vectors. The $m$th element of $\mathbf{d}$ can be expressed as
[$\mathbf{d}(\bm{\cdot})]_{m}\triangleq\tfrac{1}{\sqrt[4]{M}}\mathrm{G}\bigl{(}\tfrac{m}{\sqrt{M}}\bigr{)}\exp\bigl{(}\mathrm{j}\pi\tfrac{m}{\sqrt{M}}\bm{\cdot}\bigr{)},$ (1)
where $m\in\{-M/2,\dots,M/2-1\}$ and $\mathrm{G}:\mathbb{R}\to\mathbb{R}$ is an even, absolutely integrable function with unit $L^{2}$-norm and maximum at $\mathrm{G}(0)$. Given two random variables $\mathsf{x}$ and $\mathsf{y}$, we define a pair of transformed random vectors [^1] Without loss of generality, and for the sake of simplicity, we let $\bm{\mathsf{u}}$ and $\bm{\mathsf{v}}$ have the same dimensionality. $\bm{\mathsf{u}}\triangleq\mathbf{d}(\mathsf{x})$ and $\bm{\mathsf{v}}\triangleq\mathbf{d}(\mathsf{y})$. Their cross-covariance matrix $\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}\in\mathbb{C}^{M\times M}$ is defined as $\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}\triangleq\operatorname{\mathrm{E}}\bigl{[}\bm{\mathsf{u}}\bm{\mathsf{v}}^{\mathrm{H}}\bigr{]}-\operatorname{\mathrm{E}}[\bm{\mathsf{u}}]\operatorname{\mathrm{E}}[\bm{\mathsf{v}}]^{\mathrm{H}}$, where $\operatorname{\mathrm{E}}[\bm{\cdot}]$ denotes the expectation operator. Then, the following implication holds [5]:
$\lim_{M\to\infty}\lVert\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}\rVert_{\mathrm{F}}^{2}=0\quad\iff\quad\mathsf{x}\;{\perp\!\!\!\perp}\;\mathsf{y}\;,$ (2)
[…]
Consider $L$ i.i.d. samples of $\mathsf{x}$ and $\mathsf{y}$, $\{x(l),y(l)\}_{l=1,\dotsc,L}$. From them, we obtain the two transformed complex data matrices $\mathbf{U}\in\mathbb{C}^{M\times L}$ and $\mathbf{V}\in\mathbb{C}^{M\times L}$, defined as follows:
$\mathbf{U}$ $\triangleq[\mathbf{u}_{1},\dotsc,\mathbf{u}_{L}]=[\mathbf{d}(x(1)),\dotsc,\mathbf{d}(x(L))]$ (3a)
$\mathbf{V}$ $\triangleq[\mathbf{v}_{1},\dotsc,\mathbf{v}_{L}]=[\mathbf{d}(y(1)),\dotsc,\mathbf{d}(y(L))].$ (3b)
The unbiased sample cross-covariance between the transformed vectors is given by
$\widehat{\mathbf{C}}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}$ $\triangleq\tfrac{1}{L-1}\sum_{l=1}^{L}\Biggl{(}\mathbf{u}_{l}-\tfrac{1}{L}\sum_{i=1}^{L}\mathbf{u}_{i}\Biggr{)}\Biggl{(}\mathbf{v}_{l}-\tfrac{1}{L}\sum_{j=1}^{L}\mathbf{v}_{j}\Biggr{)}^{\mathrm{H}}$
$=\tfrac{1}{L-1}\mathbf{U}\mathbf{P}\mathbf{V}^{\mathrm{H}},\quad$ (4)
where $\mathbf{P}\triangleq\mathbf{I}-\tfrac{1}{L}\mathbf{1}\mathbf{1}^{\mathrm{T}}\in\mathbb{R}^{L\times L}$. Using (2), a marginal dependence measure is given by
$\lVert\widehat{\mathbf{C}}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}\rVert_{\mathrm{F}}^{2}=\operatorname{\mathrm{tr}}(\widehat{\mathbf{C}}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}^{\mathrm{H}}\widehat{\mathbf{C}}_{\bm{\mathsf{u}},\bm{\mathsf{v}}})$ $=\operatorname{\mathrm{tr}}\mathopen{}\mathclose{{}\left(\tfrac{1}{(L-1)^{2}}\mathbf{V}\mathbf{P}^{\mathrm{T}}\mathbf{U}^{\mathrm{H}}\mathbf{U}\mathbf{P}\mathbf{V}^{\mathrm{H}}}\right)$
$=\tfrac{1}{(L-1)^{2}}\operatorname{\mathrm{tr}}(\mathbf{P}\mathbf{U}^{\mathrm{H}}\mathbf{U}\mathbf{P}\mathbf{V}^{\mathrm{H}}\mathbf{V}),$ (5)
where the circularity of the trace operator has been used. To see the link with the HSIC, let us examine the limit for $M\to\infty$ of the Gramm (kernel) matrices $\mathbf{K}=\lim_{M\rightarrow\infty}\mathbf{U}^{\mathrm{H}}\mathbf{U}\in\mathbb{R}^{L\times L}$ and $\mathbf{Q}=\lim_{M\rightarrow\infty}\mathbf{V}^{\mathrm{H}}\mathbf{V}\in\mathbb{R}^{L\times L}$. The elements of $\mathbf{K}$ (and analogously those of $\mathbf{Q}$) are the following:
$[\mathbf{K}]_{l,l^{\prime}}$ $=\lim_{M\to\infty}\mathbf{d}^{\mathrm{H}}(x(l))\mathbf{d}(x(l^{\prime}))=$
$\lim_{M\to\infty}$ $\sum_{m=-\frac{M}{2}}^{\frac{M}{2}-1}\tfrac{1}{\sqrt{M}}\mathrm{G}^{2}\bigl{(}\tfrac{m}{\sqrt{M}}\bigr{)}\exp\biggl{(}\mathrm{j}\frac{2\pi(x(l^{\prime})-x(l))m}{\sqrt{M}}\biggr{)}$ (6a)
$=\int_{-\infty}^{\infty}\mathrm{G}^{2}(f)\exp(\mathrm{j}2\pi(x(l^{\prime})-x(l))f)\operatorname{\mathrm{d}\!}f,$ (6b)
where (6b) is the result of looking at (6a) as a Darboux sum. Then, given (II) and taking the limit $M\to\infty$, we finally obtain the HSIC [5, Sec. 3.1]:
[…]
with $\mathbf{x}=[x(1),\dotsc,x(L)]^{\mathrm{T}}$ and $\mathbf{y}=[y(1),\dotsc,y(L)]^{\mathrm{T}}$.
As an alternative to RKHS theory, the result in (7) has been derived here by means of the finite-dimensional universal mapping proposed in (1), which will play a fundamental role in the extension to the conditional case proposed later on. By defining $\kappa(\bm{\cdot})$ as the inverse Fourier transform222 From the properties imposed on $\mathrm{G}^{2}(f)$, $\kappa(\bm{\cdot})$ becomes naturally an autocorrelation translation-invariant kernel, such that $\kappa(0)=1\geq|\kappa(s)|$ and $\kappa(\infty)=\kappa(-\infty)=0$. The reader is referred to [15, Sec. 1.4] where these ideas emerge in the light of Bochner’s theorem. of $\mathrm{G}^{2}(f)$, the $(l,l^{\prime})$ entries of $\mathbf{K}$ and $\mathbf{Q}$ are just the evaluation of $\kappa(\bm{\cdot})$ at a value given by the difference between two data samples:
$[\mathbf{K}]_{l,l^{\prime}}=\kappa\bigl{(}x(l^{\prime})-x(l)\bigr{)}.$ (8)
Motivated by this structure, let us introduce a more compact and general notation that will become relevant later on
$[\mathbf{S}(\mathbf{a},\mathbf{b})]_{l,l^{\prime}}\triangleq\kappa\bigl{(}[\mathbf{b}]_{l}-[\mathbf{a}]_{l^{\prime}}\bigr{)},$ (9)
being $\mathbf{S}(\mathbf{a},\mathbf{b})$ a general kernel matrix obtained from two generic vectors $\mathbf{a}$ and $\mathbf{b}$. With it, we can compactly write both kernel matrices using (9) as follows:
$\mathbf{K}=\mathbf{S}(\mathbf{x},\mathbf{x}),\quad\mathbf{Q}=\mathbf{S}\mathopen{}\mathclose{{}\left(\mathbf{y},\mathbf{y}}\right).$ (10)
In summary, the HSIC results from measuring correlation in a finite-dimensional space based on steering vectors. From that, kernel formulations arise in a second stage once dimensionality grows without bound.
III Sample covariance matrix as an incomplete U-Statistic
Next we present the U-Statistic reformulation jointly with its incomplete expression. The objective is to provide an alternative covariance matrix estimation based on pairwise differences between the original samples that will serve as an introduction to the pruning of data pairs.
Consider a list containing all the unique $K_{\max}\triangleq L(L-1)/2$ pairs that can be constructed from $L$ different samples of $\mathsf{x}$ or $\mathsf{y}$. This list is ordered arbitrarily such that a single index $k$ identifies both elements of a pair:
$k\mapsto(x(\mathrm{f}_{1}(k)),x(\mathrm{f}_{2}(k))),\quad k\mapsto(y(\mathrm{f}_{1}(k)),y(\mathrm{f}_{2}(k))).$ (11)
Functions $\mathrm{f}_{1}(\bm{\cdot})$ and $\mathrm{f}_{2}(\bm{\cdot})$ return the corresponding indices of each term of the pair. Given $K\leq K_{\max}$ indices, we construct the pairwise differences
$\mathring{\mathbf{u}}_{k}\triangleq\frac{1}{\sqrt{2}}\mathopen{}\mathclose{{}\left(\mathbf{u}_{f_{1}(k)}-\mathbf{u}_{f_{2}(k)}}\right),\quad\mathring{\mathbf{v}}_{k}\triangleq\frac{1}{\sqrt{2}}\mathopen{}\mathclose{{}\left(\mathbf{v}_{f_{1}(k)}-\mathbf{v}_{f_{2}(k)}}\right),$ (12)
for $k\in\{1,\dots,K\}$, corresponding to the samples of the new zero-mean virtual sources $\mathring{\bm{\mathsf{u}}}$ and $\mathring{\bm{\mathsf{v}}}$. Their sample cross-covariance matrix $\widehat{\mathbf{C}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}}\in\mathbb{C}^{M\times M}$ is then the following:
$\widehat{\mathbf{C}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}}=\tfrac{1}{K}\mathring{\mathbf{U}}\mathring{\mathbf{V}}^{\mathrm{H}},$ (13)
where $\mathring{\mathbf{U}}\triangleq[\mathring{\mathbf{u}}_{1},\dots,\mathring{\mathbf{u}}_{K}]\in\mathbb{C}^{M\times K}$ and $\mathring{\mathbf{V}}\triangleq[\mathring{\mathbf{v}}_{1},\dots,\mathring{\mathbf{v}}_{K}]\in\mathbb{C}^{M\times K}$. Note that, thanks to the constant factor in (12), $\mathring{\bm{\mathsf{u}}}$ and $\mathring{\bm{\mathsf{v}}}$ have the same average covariance matrix as $\bm{\mathsf{u}}$ and $\bm{\mathsf{v}}$, respectively. In contrast to (4), $\mathbf{P}$ is missing in (13) as a result of constructing zero-mean virtual data in (12), which will lead to a cleaner implementation and fewer matrix operations.
It is worth noting that expression (13) is in fact an instance of a U-Statistic. For $K=K_{\text{max}}$, i.e. when all data pairs are taken, we obtain the equality $\widehat{\mathbf{C}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}}=\widehat{\mathbf{C}}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}$ from U-Statistics theory [11]. In contrast, for $K<K_{\text{max}}$, i.e. when (13) is an incomplete U-Statistic, although $\widehat{\mathbf{C}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}}$ remains unbiased, its estimation variance increases due to the pruning of data. Nevertheless, $\widehat{\mathbf{C}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}}$ is still a consistent estimate of $\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}$ provided that $K\rightarrow\infty$ as $L\rightarrow\infty$.
Remark 1 (Robustness to pruning)
A notable property of incomplete U-statistics is that their robustness against data pruning increases the larger $L$ is [12]. To briefly illustrate this property, consider taking only the $K=\mathopen{}\mathclose{{}\left\lfloor L/2}\right\rfloor$ data pairs with no indices in common. The number of remaining, unused, data pairs is equal to $L(L-1)/2-\lfloor L/2\rfloor$, which increases with $O(L^{2})$. To only use the i.i.d., nonrepeated, terms is effectively equivalent to computing $\widehat{\mathbf{C}}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}$ with half of the available samples [16]. It is then safe to assume that their contribution to the overall accuracy of the sample covariance is higher than those with repeated indices. Therefore, the larger $L$ is, the higher the amount of pairs that can be pruned for some specified degradation in the estimation accuracy of the resulting sample covariance. The implication is that $K$ in the incomplete U-Statistic can be designed to grow with $O(L)$ instead of $O(L^{2})$, which provides a lot of flexibility for pruning, and will be used for choosing the number of pairs in the next section.
IV Conditional dependence via U-Statistics
The conditional cross-covariance matrix between $\bm{\mathsf{u}}$ and $\bm{\mathsf{v}}$ is defined as:
$\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}|\mathsf{z}}\triangleq\int_{\mathbb{R}}\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}|\mathsf{z}=z}\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z)=\int_{\mathbb{R}}\mathbf{C}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathsf{z}=z}\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z)$ (14)
where $\mathbf{C}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathsf{z}=z}=\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}|\mathsf{z}=z}$ (given (12)) is the cross-covariance matrix conditioned to a specific value $z$ of a confounder $\mathsf{z}$, and $F_{\mathsf{z}}(z)$ is its cumulative distribution. Note that (14) is the most general definition of conditional cross-covariance that is valid for any statistics, while the Schur complement expression is only justified for Gaussian vectors [9, Ch. 4]. With the goal of deriving an estimator for (14), and following a similar rationale to $\mathring{\bm{\mathsf{u}}}$ and $\mathring{\bm{\mathsf{v}}}$, we define the virtual random variable $\mathring{\mathsf{z}}\triangleq\mathsf{z}_{1}-\mathsf{z}_{2}$, where $\mathsf{z}_{1}$ and $\mathsf{z}_{2}$ are mutually independent and distributed as $\mathsf{z}$. Given that integrating all values of $\mathsf{z}$ is equivalent to doing so for $\mathring{\mathsf{z}}=0$, i.e. $\mathsf{z}_{1}=\mathsf{z}_{2}$, the expectation in (14) can be alternatively expressed as
$\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}|\mathsf{z}}$ $=\int_{\mathbb{R}^{2}}\mathbf{C}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathring{\mathsf{z}}=0}\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z_{1})\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z_{2})$
$=\mathbf{C}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathring{\mathsf{z}}=0}\int_{\mathbb{R}^{2}}\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z_{1})\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z_{2})=\mathbf{C}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathring{\mathsf{z}}=0},$ (15)
where $\int_{\mathbb{R}^{2}}\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z_{1})\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z_{2})=1$, and since $\mathbf{C}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathring{\mathsf{z}}=0}$ does not depend on the specific values of $z_{1}$ and $z_{2}$ but rather on them being equal. In consequence, conditioning the covariance matrix with respect to $\mathsf{z}$ is equal to conditioning it with respect to $\mathring{\mathsf{z}}=0$. Therefore, the result in (IV) suggests letting the potential confounder data control the pruning of the incomplete U-Statistics in (13). But, since $\mathring{\mathsf{z}}=0$ is an event of zero probability for continuous random variables, the data control should be based on merely small values of $|\mathring{z}|$. Note that the procedure followed here is the same followed in [12].
With the intention of choosing the data pairs for the pruning based on $|\mathring{z}|$, we define the samples of $\mathring{z}$ as follows:
$\mathring{z}_{l,l^{\prime}}\triangleq z(l)-z(l^{\prime}),$ (16)
where $z(l)$ and $z(l^{\prime})$ are i.i.d. samples drawn from $\mathsf{z}$ with $l\neq l^{\prime}$. Then, we let the sorting of $|\mathring{z}_{l,l^{\prime}}|$ (in ascending order) be the one that determines the ordering of the index pairs provided by $\mathrm{f}_{1}(\bm{\cdot})$ and $\mathrm{f}_{2}(\bm{\cdot})$ in (11). Moreover, in view of Remark 1, the amount of pairs $K$ is set to grow as $O(L)$, i.e.
$K_{\alpha}\triangleq\bigl{\lfloor}\tfrac{L\alpha}{2}\bigr{\rfloor},$ (17)
being $1\leq\alpha\ll(L-1)$ a tuning hyper-parameter. While $\alpha=1$ ensures that only very small values of $|\mathring{z}_{l,l^{\prime}}|$ are considered, in the other extreme of $\alpha=L-1$ the U-Statistic becomes complete and there is no conditioning at all, thus encountering the HSIC as a particular case as in (7). A natural trade-off then emerges: while $\alpha$ close to $1$ is desirable for the conditioning idea (IV) to work properly, higher values of $\alpha$ will avoid excessive pruning and provide sufficient statistical accuracy in (13) with $K=K_{\alpha}$. Remarkably, the selection of the hyper-parameter $\alpha$ becomes a minor issue provided that $L$ is sufficiently large, as was shown in [12] under the correlation measure framework between a pair of vectors, and will be seen later on by a numerical example.
IV-A Conditional HSIC
Now that we have determined the sorting and pruning of the data pairs according to the confounder $\mathsf{z}$, let us write a conditional dependence measure as the Frobenius norm of (13):
$\operatorname{\mathrm{tr}}\bigl{(}\widehat{\mathbf{C}}^{\mathrm{H}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathsf{z}}\widehat{\mathbf{C}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathsf{z}}\bigr{)}=\tfrac{1}{K_{\alpha}^{2}}\operatorname{\mathrm{tr}}\bigl{(}\mathring{\mathbf{U}}^{\mathrm{H}}\mathring{\mathbf{U}}\mathring{\mathbf{V}}^{\mathrm{H}}\mathring{\mathbf{V}}\bigr{)},$ (18)
where the circularity of the trace has been used. To link the previous expression with kernel-based methods, let us rewrite the zero-mean virtual data matrices $\mathring{\mathbf{U}}$ and $\mathring{\mathbf{V}}$ as follows:
$\mathring{\mathbf{U}}=\tfrac{1}{\sqrt{2}}(\mathbf{U}_{1}-\mathbf{U}_{2}),\qquad\mathring{\mathbf{V}}=\tfrac{1}{\sqrt{2}}(\mathbf{V}_{1}-\mathbf{V}_{2}),$ (19)
where
$\mathbf{U}_{a}\triangleq[\mathbf{u}_{\mathrm{f}_{a}(1)},\dots,\mathbf{u}_{\mathrm{f}_{a}(K_{\alpha})}],\;\;\mathbf{V}_{a}\triangleq[\mathbf{v}_{\mathrm{f}_{a}(1)},\dots,\mathbf{v}_{\mathrm{f}_{a}(K_{\alpha})}],$ (20)
for $a=\{1,2\}$. Accordingly, (18) is then rewritten as
$\operatorname{\mathrm{tr}}\bigl{(}\widehat{\mathbf{C}}^{\mathrm{H}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathsf{z}}\widehat{\mathbf{C}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathsf{z}}\bigr{)}$ (21)
$=\tfrac{1}{4K_{\alpha}^{2}}\operatorname{\mathrm{tr}}\bigl{(}(\mathbf{U}_{1}-\mathbf{U}_{2})^{\mathrm{H}}(\mathbf{U}_{1}-\mathbf{U}_{2})(\mathbf{V}_{1}-\mathbf{V}_{2})^{\mathrm{H}}(\mathbf{V}_{1}-\mathbf{V}_{2})\bigr{)}.$
Taking the limit $M\rightarrow\infty$, kernel matrices are obtained from the products among $\mathbf{U}$ and $\mathbf{V}$ in (21):
$\mathbf{K}_{a,a^{\prime}}\triangleq\lim_{M\rightarrow\infty}\mathbf{U}_{a}^{\mathrm{H}}\mathbf{U}_{a^{\prime}}$ $=\mathbf{S}\mathopen{}\mathclose{{}\left(\mathbf{x}_{a},\mathbf{x}_{a^{\prime}}}\right)\quad$ (22a)
$\mathbf{Q}_{a,a^{\prime}}\triangleq\lim_{M\rightarrow\infty}\mathbf{V}_{a}^{\mathrm{H}}\mathbf{V}_{a^{\prime}}$ $=\mathbf{S}\mathopen{}\mathclose{{}\left(\mathbf{y}_{a},\mathbf{y}_{a^{\prime}}}\right),$ (22b)
where $\mathbf{S}\mathopen{}\mathclose{{}\left(\bm{\cdot},\bm{\cdot}}\right)$ is given in (9). Therefore, the new relevant data subsets for inferring conditional dependence are given by
$[\mathbf{x}_{a}]_{k}\triangleq x(f_{a}(k)),\quad[\mathbf{y}_{a}]_{k}\triangleq y(f_{a}(k)).$ (23)
Finally, the resulting C-HSIC can be expressed as follows:
[…]
with $\breve{\mathbf{K}}\triangleq\mathbf{K}_{1,1}+\mathbf{K}_{2,2}-\mathbf{K}_{1,2}-\mathbf{K}_{2,1}$ and $\breve{\mathbf{Q}}\triangleq\mathbf{Q}_{1,1}+\mathbf{Q}_{2,2}-\mathbf{Q}_{1,2}-\mathbf{Q}_{2,1}$. Note that, as a result of the U-Statistics implementation, each entry of the new kernel-based matrices involves four data samples of the same source ($4$-tuples), in contrast to only the pairs typically involved in classical kernel methods. This fact, along with the lack of $\mathbf{P}$, are the main distinctive features of the C-HSIC (24) vs. HSIC (7).
V Numerical illustration
In order to test the proposed method, we aim at generating uncorrelated data with a controlled amount of co-information [17] (also called interaction information [18]), defined as
$\mathrm{I}(\mathsf{x};\mathsf{y};\mathsf{z})=\mathrm{I}(\mathsf{x};\mathsf{y})-\mathrm{I}(\mathsf{x};\mathsf{y}|\mathsf{z}),$ (25)
such that the potentials of the method can be clearly highlighted. Similar ideas on modeling conditional dependences can be found in [19]. Since co-information can be either positive or negative, two scenarios are studied: $\mathcal{M}^{+}$ for $\mathrm{I}(\mathsf{x};\mathsf{y};\mathsf{z})>0$ and $\mathcal{M}^{-}$ for $\mathrm{I}(\mathsf{x};\mathsf{y};\mathsf{z})<0$, whose variables are as follows:
+:$\mathcal{M}^{+}:\begin{cases}\mathsf{x}=\sqrt{\gamma}\mathsf{a}\mathsf{p}+\mathsf{v}\\ \mathsf{y}=\sqrt{\gamma}\mathsf{a}\mathsf{q}+\mathsf{w}\\ \mathsf{z}=\mathsf{a}\end{cases},\hskip 5.0pt\mathcal{M}^{-}:\begin{cases}\mathsf{x}=\sqrt{\gamma}\mathsf{b}\mathsf{p}+\mathsf{v}\\ \mathsf{y}=\sqrt{\gamma}\mathsf{c}\mathsf{q}+\mathsf{w}\\ \mathsf{z}=\mathsf{b}-\mathsf{c}\end{cases}.$ (26)
The internal i.i.d. random variables are distributed as $\mathsf{a},\mathsf{b},\mathsf{c}\sim\mathcal{U}(0,\sqrt{3\text{}})$ (uniform), $\mathsf{v},\mathsf{w}\sim\mathcal{N}(0,1)$ (normal), and $\mathsf{p},\mathsf{q}\sim\mathrm{Bern}_{1/2}\{-1,1\}$ (equiprobable Bernoulli). Parameter $\gamma$ is the signal-to-noise ratio associated to the measurements and controls the total amount of absolute co-information. In model $\mathcal{M}^{+}$, $\mathsf{x}$ and $\mathsf{y}$ are dependent due to the influence of $\mathsf{a}$ in both, but they are conditionally independent, since knowing $\mathsf{z}$ (i.e. $\mathsf{a}$) implies that $\mathsf{x}$ and $\mathsf{y}$ become solely driven by independent phenomena ($\mathsf{v}$ and $\mathsf{w}$). By contrast, $\mathsf{x}$ and $\mathsf{y}$ are marginally independent in model $\mathcal{M}^{-}$, but they become conditionally dependent, since the knowledge of $\mathsf{z}$ correlates the possible joint values of $\mathsf{b}$ and $\mathsf{c}$. In both models, $\mathsf{x}$, $\mathsf{y}$ and $\mathsf{z}$ are mutually uncorrelated due to the multiplicative effect of mutually independent variables $\mathsf{p}$ and $\mathsf{q}$, so correlation measures are unable to discover any data association.
Regarding the choice of kernel, the universal Gaussian kernel is used, which yields $\kappa(s)=\exp(-(\frac{s}{\hat{\sigma}L^{-1/5}})^{2})$. Note that the kernel length scales in proportion with the sample standard deviation $\hat{\sigma}$ of the data and in inverse proportion with the data size $L$. The reader is referred to [7, Appendix D][20, 21] for a justification of the typical power law $O(L^{-1/5})$.
Figure 1: Model $\mathcal{M}^{+}$. Orange: $L=100$; blue: $L=600$. Markers denote the empirical average while bands indicate the standard deviation. (a)-cross marker: HSIC; (a)-circle marker: C-HSIC; (b): C-HSIC with random pruning.
Figure 2: Model $\mathcal{M}^{-}$. Orange: $L=100$; blue: $L=600$. Markers denote the empirical average while bands indicate the standard deviation. (a)-cross marker: HSIC; (a)-circle marker: C-HSIC; (b): C-HSIC with random pruning.
In Fig. 1 and Fig. 2, the measures of dependence are shown as a function of $\gamma$ for both models with a moderate hyper-parameter $\alpha=4$ from (17). This yields $8\%$ of the total data pairs for $L=100$ and $1.3\%$ for $L=600$ (and even less for higher $L$). For model $\mathcal{M}^{+}$ (Fig. 1), the C-HSIC correctly depicts small conditional dependence, while the HSIC confirms that marginal dependence is high for moderate $\gamma$ values. Conversely, Fig. 2 exhibits the capability of the C-HSIC to discover significant conditional dependence for moderate $\gamma$ values for model $\mathcal{M}^{-}$, while marginal dependence is confirmed to be small by the HSIC at any value of $\gamma$.
To get further insights on the proposed ideas, Fig. 1-(b) and Fig. 2-(b) show the performance of the C-HSIC when the pruning of data pairs in the U-Statistics is random, which means that the pair $(l(k),l^{\prime}(k))$ is not controlled by the confounder. It is seen that the C-HSIC produces a measure that mimics the marginal dependence one, thus confirming the main property of unbiasedness related to the general U-Statistics theory, and also showing its consequent increase in variance. It is also important to note that the selected value of $\alpha$, and the small percentage of data that yields this choice, confirms that the increase in variance induced by the pruning becomes small for sufficiently large data sizes due to Remark 1, making $\alpha$ a noncritical free hyper-parameter.
VI Conclusions
The classical HSIC measure for marginal statistical dependence has been reinterpreted as a correlation measure on a finite but high-dimensional space based on windowed steering vectors. From the original mapping, we can establish an insightful connection with kernels by letting the dimension grow to infinity. Instead of being merely a reinterpretation of the HSIC, this vision demonstrates its main power under the more involved conditional dependence framework. In particular, the illustrated identification of the HSIC as a sample covariance estimator opens the possibility of leveraging U-Statistics for this task. Thanks to this formulation, we can provide a novel route for performing statistical conditioning by pruning data pairs based on the pairwise differences of the confounder under suspicion. Furthermore, the proposed measure of conditional dependence does not require matrix inversions, which has the advantage of reduced computational complexity and the avoidance of addressing ill-conditioned matrices. Further work should study the potential of the proposed method with richer data sets.

### v2

@@
 Abstract
-The problem of measuring conditional dependence between two random phenomena arises when a third one (a confounder) has a potential influence on the amount of information shared by the original pair.
+The problem of measuring conditional dependence between two random phenomena arises when a third one (a confounder) has a potential influence on the amount of information between them.
 A typical issue in this challenging problem is the inversion of ill-conditioned autocorrelation matrices.
 This paper presents a novel measure of conditional dependence based on the use of incomplete unbiased statistics of degree two, which allows to re-interpret independence as uncorrelatedness on a finite-dimensional feature space.
-This formulation enables to prune data according to the observations of the confounder itself, thus avoiding matrix inversions altogether.
-Moreover, the proposed approach is articulated as an extension of the Hilbert-Schmidt independence criterion, which becomes expressible through kernels that operate on 4-tuples of data.
+This formulation enables to prune data according to observations of the confounder itself, thus avoiding matrix inversions altogether.
+The proposed approach is articulated as an extension of the Hilbert-Schmidt independence criterion, which becomes expressible through kernels that operate on 4-tuples of data.
 Index Terms:
-Hilbert-Schmidt independence criterion (HSIC), kernel methods, co-information, U-Statistics, conditional dependence.
+Hilbert-Schmidt independence criterion (HSIC), kernel methods, conditional dependence, U-Statistics.
 I Introduction
-The problem of measuring statistical dependence emerges in many fields, such as multivariate statistical analysis [1, Ch.
-6], wireless sensor networks [2, Ch.
-2], data analytics [3], and statistical array processing [4].
-The Hilbert-Schmidt independence criterion (HSIC[5] provides a versatile non-negative measure of statistical dependence that becomes null if and only if the data pair is independent.
-Grounded in the theory of reproducing kernel Hilbert space s (RKHS), the HSIC succeeds in discovering both linear and nonlinear relationships in data.
-A well-known interpretation of the HSIC is that dependence produces a correlation of distances [6], i.e.
-close couples in one data set coincide with close couples in the other data set.
-Another insightful interpretation is that the HSIC measures statistical dependence by gauging the correlation of data implicitly mapped onto an infinite-dimensional space.
-This basic idea of translating dependence to correlation also admits a more traditional implementation by explicitly mapping data onto a finite-dimensional space [7], with the advantage of leveraging classical second-order statistics for inferring information-theoretic measures.
-Conditional dependence becomes relevant when a third phenomenon might explain, mediate, or confound the apparent relationship exhibited by the original random pair, which happens in fields such as causal discovery and Bayesian network learning [8].
-Although mapping data onto a space of higher dimension is still useful for that purpose, it is still necessary to measure conditional correlation in this newfound space.
-This method usually leads to the computation of the Schur complement of high-dimensional matrices [9, Ch.
-4], which is known to have numerical issues concerned with matrix inversions and regularization due to the usually low rank structure of the involved autocorrelation matrices [10].
-This letter presents a procedure for measuring conditional dependence by statistically conditioning the observed data to a potential confounder, which bypasses the matrix inversion problems.
-First, a novel theoretical derivation of the classical HSIC is provided in Sec.
-II, by translating the problem of statistical dependence into one of correlation after mapping the data onto finite-dimensional spaces based on steering vectors [7].
+Conditional dependence becomes relevant when a third phenomenon might explain, mediate, or confound the apparent relationship exhibited by an original random pair.
+Its measurement is an important task in causal discovery and Bayesian network learning [1, 2, 3, Ch.
+7], which emerge in applied fields such as earth system sciences [4] or clinical data analysis [5].
+Due to its nature, empirically estimating conditional dependence is a complex problem, given that it also requires inferring conditional marginal distributions.
+A known approach consists in measuring the Hilbert-Schmidt norm of the normalized conditional cross-covariance operator [6], grounded in the theory of reproducing kernel Hilbert space s (RKHS).
+This kernel-based method is capable of measuring statistical dependence by gauging correlation of data implicitly mapped onto an infinite-dimensional space.
+However, it is also known to have numerical issues concerned with matrix inversions and regularization due to the usually low-rank structure of the involved correlation matrices [6].
+This also leads to a fairly complex method that scales cubically with the observed data size, becoming prohibitive for large data sets.
+These issues have lately been the subject of interest of numerous lines of research.
+A kernel-based conditional independence test is proposed in [1] by deriving the asymptotic distribution of the appropriate test statistic under the null hypothesis, which succeeds in avoiding matrix inversions, but whose complexity still increases cubically with the sample size.
+Other authors have proposed to approximate kernel functions by randomly selecting features [7], by providing an alternative review to the embedding process onto the RKHS [8], or by leveraging the interpretation that dependence produces correlation of distances [9], i.e.
+close pairs in one data set coincide with close pairs in the other [10].
+Nevertheless, none of these methods addresses the aforementioned computational issues, since they all require matrix inversions.
+An extension of the Hilbert-Schmidt independence criterion (HSIC[11], a measure of unconditional independence, is introduced in [12] with the objective of inferring causality, a similar endeavor to this letter, although focused on joint independence rather than conditional dependence.
+This letter presents a procedure for measuring conditional dependence by statistically conditioning the observed data to a potential confounder in a simple way, bypassing matrix inversion problems.
+The basis for conditioning is to prune pairwise differences of data under the control of the confounder [13], stemming from distance-based methods [10].
+Simultaneously, the estimate is constructed as an alternative derivation of the kernel-based HSIC.
+While these two approaches are known to be equivalent [14], we propose here an estimate that combines them, albeit for different tasks.
+To do so, we first provide a novel reinterpretation of the HSIC in Sec.
+II by translating the problem of statistical dependence into one of correlation after mapping data onto finite-dimensional spaces based on steering vectors [15].
 Sec.
 III briefly reviews the theory of unbiased statistics (U-Statistics) [16], which is then employed in Sec.
-IV to show that conditional dependence can be accomplished by just pruning U-Statistics under the control of the confounder [12].
-The obtained measure, the conditional HSIC (C-HSIC), is based on a signal processing structure resembling kernel methods, which embraces the HSIC as a particular case when the U-Statistic is complete.
-However, it has the distinctive feature of being articulated on 4-tuples instead of the classical processing based on pairs of data from each phenomenon.
+IV to show that conditional dependence can be obtained by pruning U-Statistics.
+The obtained measure, named conditional HSIC (C-HSIC), is based on a classical signal processing structure but linked to kernel methods, which embraces the HSIC as a particular case when the U-Statistic is complete.
 II Marginal dependence as correlation
 Let $\mathbf{d}:\mathbb{R}\to\mathbb{C}^{M}$ be a mapping based on windowed steering vectors. The $m$th element of $\mathbf{d}$ can be expressed as
-[$\mathbf{d}(\bm{\cdot})]_{m}\triangleq\tfrac{1}{\sqrt[4]{M}}\mathrm{G}\bigl{(}\tfrac{m}{\sqrt{M}}\bigr{)}\exp\bigl{(}\mathrm{j}\pi\tfrac{m}{\sqrt{M}}\bm{\cdot}\bigr{)},$ (1)
-where $m\in\{-M/2,\dots,M/2-1\}$ and $\mathrm{G}:\mathbb{R}\to\mathbb{R}$ is an even, absolutely integrable function with unit $L^{2}$-norm and maximum at $\mathrm{G}(0)$.
-Given two random variables $\mathsf{x}$ and $\mathsf{y}$, we define a pair of transformed random vectors [^1] Without loss of generality, and for the sake of simplicity, we let $\bm{\mathsf{u}}$ and $\bm{\mathsf{v}}$ have the same dimensionality.
-$\bm{\mathsf{u}}\triangleq\mathbf{d}(\mathsf{x})$ and $\bm{\mathsf{v}}\triangleq\mathbf{d}(\mathsf{y})$.
-Their cross-covariance matrix $\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}\in\mathbb{C}^{M\times M}$ is defined as $\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}\triangleq\operatorname{\mathrm{E}}\bigl{[}\bm{\mathsf{u}}\bm{\mathsf{v}}^{\mathrm{H}}\bigr{]}-\operatorname{\mathrm{E}}[\bm{\mathsf{u}}]\operatorname{\mathrm{E}}[\bm{\mathsf{v}}]^{\mathrm{H}}$, where $\operatorname{\mathrm{E}}[\bm{\cdot}]$ denotes the expectation operator.
-Then, the following implication holds [5]:
-$\lim_{M\to\infty}\lVert\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}\rVert_{\mathrm{F}}^{2}=0\quad\iff\quad\mathsf{x}\;{\perp\!\!\!\perp}\;\mathsf{y}\;,$ (2)
+[$\mathbf{d}(\bm{\cdot})]_{m}\triangleq\tfrac{1}{\sqrt[4]{M}}\mathrm{G}\bigl{(}\tfrac{m}{\sqrt{M}}\bigr{)}\mathit{e}^{\mathrm{j}\pi\frac{m}{\sqrt{M}}\bm{\cdot}},$ (1)
 […]
+$\lim_{M\to\infty}\lVert\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}\rVert_{\mathrm{F}}^{2}=0\quad\iff\quad\mathsf{x},\mathsf{y}\text{ are independent},$ (2)
+where $\lVert\bm{\cdot}\rVert_{\mathrm{F}}$ denotes the Frobenius norm.
+Terms $\operatorname{\mathrm{E}}[\mathbf{d}(\mathsf{x})]$ and $\operatorname{\mathrm{E}}[\mathbf{d}(\mathsf{y})]$ become dense samplings of the weighted (by $\mathrm{G}(\bm{\cdot})$) marginal characteristic function of $\mathsf{x}$ and $\mathsf{y}$, respectively, in the limit of $M\to\infty$.
+Similarly, $\operatorname{\mathrm{E}}[\mathbf{d}(\mathsf{x})\mathbf{d}^{\mathrm{H}}(\mathsf{y})]$ becomes a dense sampling of the weighted joint characteristic function (JCF).
+Therefore, (2) is effectively checking the separability of the JCF at every sample point (i.e.
+the factorization of $\operatorname{\mathrm{E}}[\mathbf{d}(\mathsf{x})\mathbf{d}^{\mathrm{H}}(\mathsf{y})]$ as a product of expectations), which becomes equivalent to an uncorrelatedness property [17, 15], akin to the distance covariance in [10].
 Consider $L$ i.i.d.
-samples of $\mathsf{x}$ and $\mathsf{y}$, $\{x(l),y(l)\}_{l=1,\dotsc,L}$.
-From them, we obtain the two transformed complex data matrices $\mathbf{U}\in\mathbb{C}^{M\times L}$ and $\mathbf{V}\in\mathbb{C}^{M\times L}$, defined as follows:
+samples of $\mathsf{x}$ and $\mathsf{y}$, $\{x(l),y(l)\}_{l=1,\dotsc,L}$, from which we obtain two transformed complex data matrices $\mathbf{U}\in\mathbb{C}^{M\times L}$ and $\mathbf{V}\in\mathbb{C}^{M\times L}$, defined as follows:
 $\mathbf{U}$ $\triangleq[\mathbf{u}_{1},\dotsc,\mathbf{u}_{L}]=[\mathbf{d}(x(1)),\dotsc,\mathbf{d}(x(L))]$ (3a)
 $\mathbf{V}$ $\triangleq[\mathbf{v}_{1},\dotsc,\mathbf{v}_{L}]=[\mathbf{d}(y(1)),\dotsc,\mathbf{d}(y(L))].$ (3b)
-The unbiased sample cross-covariance between the transformed vectors is given by
+The sample cross-covariance between the transformed vectors is given by
 $\widehat{\mathbf{C}}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}$ $\triangleq\tfrac{1}{L-1}\sum_{l=1}^{L}\Biggl{(}\mathbf{u}_{l}-\tfrac{1}{L}\sum_{i=1}^{L}\mathbf{u}_{i}\Biggr{)}\Biggl{(}\mathbf{v}_{l}-\tfrac{1}{L}\sum_{j=1}^{L}\mathbf{v}_{j}\Biggr{)}^{\mathrm{H}}$
 $=\tfrac{1}{L-1}\mathbf{U}\mathbf{P}\mathbf{V}^{\mathrm{H}},\quad$ (4)
-where $\mathbf{P}\triangleq\mathbf{I}-\tfrac{1}{L}\mathbf{1}\mathbf{1}^{\mathrm{T}}\in\mathbb{R}^{L\times L}$. Using (2), a marginal dependence measure is given by
-$\lVert\widehat{\mathbf{C}}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}\rVert_{\mathrm{F}}^{2}=\operatorname{\mathrm{tr}}(\widehat{\mathbf{C}}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}^{\mathrm{H}}\widehat{\mathbf{C}}_{\bm{\mathsf{u}},\bm{\mathsf{v}}})$ $=\operatorname{\mathrm{tr}}\mathopen{}\mathclose{{}\left(\tfrac{1}{(L-1)^{2}}\mathbf{V}\mathbf{P}^{\mathrm{T}}\mathbf{U}^{\mathrm{H}}\mathbf{U}\mathbf{P}\mathbf{V}^{\mathrm{H}}}\right)$
+where $\mathbf{P}\triangleq\mathbf{I}-\tfrac{1}{L}\mathbf{1}\mathbf{1}^{\mathrm{T}}\in\mathbb{R}^{L\times L}$ is the data centering matrix, a well-known projection matrix in the signal processing and kernel methods literature [18, Appx.
+B.7].
+Using (2), a marginal dependence measure is given by
+$\lVert\widehat{\mathbf{C}}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}\rVert_{\mathrm{F}}^{2}=\operatorname{\mathrm{tr}}(\widehat{\mathbf{C}}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}^{\mathrm{H}}\widehat{\mathbf{C}}_{\bm{\mathsf{u}},\bm{\mathsf{v}}})$ $=\operatorname{\mathrm{tr}}\mathopen{}\mathclose{{}\left(\tfrac{1}{(L-1)^{2}}\mathbf{V}\mathbf{P}^{\mathrm{H}}\mathbf{U}^{\mathrm{H}}\mathbf{U}\mathbf{P}\mathbf{V}^{\mathrm{H}}}\right)$
 $=\tfrac{1}{(L-1)^{2}}\operatorname{\mathrm{tr}}(\mathbf{P}\mathbf{U}^{\mathrm{H}}\mathbf{U}\mathbf{P}\mathbf{V}^{\mathrm{H}}\mathbf{V}),$ (5)
 where the circularity of the trace operator has been used.
-To see the link with the HSIC, let us examine the limit for $M\to\infty$ of the Gramm (kernel) matrices $\mathbf{K}=\lim_{M\rightarrow\infty}\mathbf{U}^{\mathrm{H}}\mathbf{U}\in\mathbb{R}^{L\times L}$ and $\mathbf{Q}=\lim_{M\rightarrow\infty}\mathbf{V}^{\mathrm{H}}\mathbf{V}\in\mathbb{R}^{L\times L}$.
+To see the link of (II) with the HSIC, let us examine the limit for $M\to\infty$ of the kernel matrices $\mathbf{K}\triangleq\lim_{M\rightarrow\infty}\mathbf{U}^{\mathrm{H}}\mathbf{U}\in\mathbb{R}^{L\times L}$ and $\mathbf{Q}\triangleq\lim_{M\rightarrow\infty}\mathbf{V}^{\mathrm{H}}\mathbf{V}\in\mathbb{R}^{L\times L}$.
 The elements of $\mathbf{K}$ (and analogously those of $\mathbf{Q}$) are the following:
-$[\mathbf{K}]_{l,l^{\prime}}$ $=\lim_{M\to\infty}\mathbf{d}^{\mathrm{H}}(x(l))\mathbf{d}(x(l^{\prime}))=$
-$\lim_{M\to\infty}$ $\sum_{m=-\frac{M}{2}}^{\frac{M}{2}-1}\tfrac{1}{\sqrt{M}}\mathrm{G}^{2}\bigl{(}\tfrac{m}{\sqrt{M}}\bigr{)}\exp\biggl{(}\mathrm{j}\frac{2\pi(x(l^{\prime})-x(l))m}{\sqrt{M}}\biggr{)}$ (6a)
-$=\int_{-\infty}^{\infty}\mathrm{G}^{2}(f)\exp(\mathrm{j}2\pi(x(l^{\prime})-x(l))f)\operatorname{\mathrm{d}\!}f,$ (6b)
-where (6b) is the result of looking at (6a) as a Darboux sum. Then, given (II) and taking the limit $M\to\infty$, we finally obtain the HSIC [5, Sec. 3.1]:
+$[\mathbf{K}]_{l,l^{\prime}}$ =${\lim_{M\to\infty}^{}}\mathbf{d}^{\mathrm{H}}(x(l))\mathbf{d}(x(l^{\prime}))$
+=${\lim_{M\to\infty}^{}}\smashoperator[]{\sum_{m=-\frac{M}{2}}^{\frac{M}{2}-1}}\tfrac{1}{\sqrt{M}}\mathrm{G}^{2}\bigl{(}\tfrac{m}{\sqrt{M}}\bigr{)}\mathit{e}^{-\mathrm{j}2\pi(x(l)-x(l^{\prime}))\frac{m}{\sqrt{M}}}$ (6)
+$=\int_{\mathrlap{-\infty}}^{\mathrlap{\infty}}\mathrm{G}^{2}(f)\mathit{e}^{-\mathrm{j}2\pi(x(l)-x(l^{\prime}))f}\operatorname{\mathrm{d}\!}f\triangleq\kappa\bigl{(}x(l)-x(l^{\prime})\bigr{)},$
+where $\kappa(\bm{\cdot})$ is the kernel function that results from the Fourier transform333 From the properties imposed on $\mathrm{G}^{2}(f)$, $\kappa(\bm{\cdot})$ becomes naturally an autocorrelation translation-invariant kernel, such that $\kappa(0)=1\geq\lvert\kappa(s)\rvert$ and $\kappa(\infty)=\kappa(-\infty)=0$.
+The reader is referred to [19, Sec.
+1.4] where these ideas emerge in the light of Bochner’s theorem.
+of $\mathrm{G}^{2}(f)$, and the integral is the limit of the Darboux sum in the second line.
+As a result, the entries of $\mathbf{K}$ and $\mathbf{Q}$ are just the evaluation of $\kappa(\bm{\cdot})$ at the difference between two data samples.
+Then, given (II) and taking the limit $M\to\infty$, we obtain the HSIC [11, Sec.
+3.1]:
 […]
 with $\mathbf{x}=[x(1),\dotsc,x(L)]^{\mathrm{T}}$ and $\mathbf{y}=[y(1),\dotsc,y(L)]^{\mathrm{T}}$.
-As an alternative to RKHS theory, the result in (7) has been derived here by means of the finite-dimensional universal mapping proposed in (1), which will play a fundamental role in the extension to the conditional case proposed later on.
-By defining $\kappa(\bm{\cdot})$ as the inverse Fourier transform222 From the properties imposed on $\mathrm{G}^{2}(f)$, $\kappa(\bm{\cdot})$ becomes naturally an autocorrelation translation-invariant kernel, such that $\kappa(0)=1\geq|\kappa(s)|$ and $\kappa(\infty)=\kappa(-\infty)=0$.
-The reader is referred to [15, Sec.
-1.4] where these ideas emerge in the light of Bochner’s theorem.
-of $\mathrm{G}^{2}(f)$, the $(l,l^{\prime})$ entries of $\mathbf{K}$ and $\mathbf{Q}$ are just the evaluation of $\kappa(\bm{\cdot})$ at a value given by the difference between two data samples:
-$[\mathbf{K}]_{l,l^{\prime}}=\kappa\bigl{(}x(l^{\prime})-x(l)\bigr{)}.$ (8)
-Motivated by this structure, let us introduce a more compact and general notation that will become relevant later on
-$[\mathbf{S}(\mathbf{a},\mathbf{b})]_{l,l^{\prime}}\triangleq\kappa\bigl{(}[\mathbf{b}]_{l}-[\mathbf{a}]_{l^{\prime}}\bigr{)},$ (9)
-being $\mathbf{S}(\mathbf{a},\mathbf{b})$ a general kernel matrix obtained from two generic vectors $\mathbf{a}$ and $\mathbf{b}$.
-With it, we can compactly write both kernel matrices using (9) as follows:
-$\mathbf{K}=\mathbf{S}(\mathbf{x},\mathbf{x}),\quad\mathbf{Q}=\mathbf{S}\mathopen{}\mathclose{{}\left(\mathbf{y},\mathbf{y}}\right).$ (10)
-In summary, the HSIC results from measuring correlation in a finite-dimensional space based on steering vectors.
-From that, kernel formulations arise in a second stage once dimensionality grows without bound.
+In summary, the alternative formulation exposed above shows that the HSIC results from measuring correlation in a finite-dimensional space based on steering vectors, where the kernel formulation arises in a second stage once the dimensionality grows without bound.
+The interplay between kernels and cross-covariance matrices of mapped data will be used later on to condition the kernel matrices by pruning data.
+Beforehand, we will reformulate the estimation of cross-covariance matrices leveraging incomplete U-statistics, which will serve as the basis for data pruning.
 III Sample covariance matrix as an incomplete U-Statistic
-Next we present the U-Statistic reformulation jointly with its incomplete expression.
-The objective is to provide an alternative covariance matrix estimation based on pairwise differences between the original samples that will serve as an introduction to the pruning of data pairs.
-Consider a list containing all the unique $K_{\max}\triangleq L(L-1)/2$ pairs that can be constructed from $L$ different samples of $\mathsf{x}$ or $\mathsf{y}$.
-This list is ordered arbitrarily such that a single index $k$ identifies both elements of a pair:
-$k\mapsto(x(\mathrm{f}_{1}(k)),x(\mathrm{f}_{2}(k))),\quad k\mapsto(y(\mathrm{f}_{1}(k)),y(\mathrm{f}_{2}(k))).$ (11)
-Functions $\mathrm{f}_{1}(\bm{\cdot})$ and $\mathrm{f}_{2}(\bm{\cdot})$ return the corresponding indices of each term of the pair. Given $K\leq K_{\max}$ indices, we construct the pairwise differences
-$\mathring{\mathbf{u}}_{k}\triangleq\frac{1}{\sqrt{2}}\mathopen{}\mathclose{{}\left(\mathbf{u}_{f_{1}(k)}-\mathbf{u}_{f_{2}(k)}}\right),\quad\mathring{\mathbf{v}}_{k}\triangleq\frac{1}{\sqrt{2}}\mathopen{}\mathclose{{}\left(\mathbf{v}_{f_{1}(k)}-\mathbf{v}_{f_{2}(k)}}\right),$ (12)
-for $k\in\{1,\dots,K\}$, corresponding to the samples of the new zero-mean virtual sources $\mathring{\bm{\mathsf{u}}}$ and $\mathring{\bm{\mathsf{v}}}$.
+Consider a list containing all the unique $K_{\max}\triangleq L(L-1)/2$ pairs that can be constructed from $L$ different samples of a random variable.
+This list is ordered arbitrarily such that a single index $k$ identifies both elements of a pair.
+Let $\mathrm{f}_{1}(k)$ and $\mathrm{f}_{2}(k)$ be two functions that return the corresponding indices of each term of the $k$th pair.
+Given $K\leq K_{\max}$ indices, we construct the pairwise differences
+$\mathring{\mathbf{u}}_{k}\triangleq\tfrac{1}{\sqrt{2}}\mathopen{}\mathclose{{}\left(\mathbf{u}_{\mathrm{f}_{1}(k)}-\mathbf{u}_{\mathrm{f}_{2}(k)}}\right),\quad\mathring{\mathbf{v}}_{k}\triangleq\tfrac{1}{\sqrt{2}}\mathopen{}\mathclose{{}\left(\mathbf{v}_{\mathrm{f}_{1}(k)}-\mathbf{v}_{\mathrm{f}_{2}(k)}}\right),$ (8)
+for $k\in\{1,\dots,K\}$, corresponding to the samples of two virtual sources $\mathring{\bm{\mathsf{u}}}$ and $\mathring{\bm{\mathsf{v}}}$.
+It is worth noting that these are zero-mean variables for any pair with $\mathrm{f}_{1}(k)\neq\mathrm{f}_{2}(k)$, since $\bm{\mathsf{u}}$ and $\bm{\mathsf{v}}$ are i.i.d..
 Their sample cross-covariance matrix $\widehat{\mathbf{C}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}}\in\mathbb{C}^{M\times M}$ is then the following:
@@
 In contrast to (4), $\mathbf{P}$ is missing in (9) as a result of constructing zero-mean virtual data in (8), which will lead to a cleaner implementation and fewer matrix operations.
-It is worth noting that expression (13) is in fact an instance of a U-Statistic.
+Equation (9) is in fact an instance of a U-Statistic.
 For $K=K_{\text{max}}$, i.e.
@@
 The number of remaining, unused, data pairs is equal to $L(L-1)/2-\lfloor L/2\rfloor$, which increases with $O(L^{2})$.
-To only use the i.i.d., nonrepeated, terms is effectively equivalent to computing $\widehat{\mathbf{C}}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}$ with half of the available samples [16].
+To only use the i.i.d., unique, terms is effectively equivalent to computing $\widehat{\mathbf{C}}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}$ with half of the available samples [20].
 It is then safe to assume that their contribution to the overall accuracy of the sample covariance is higher than those with repeated indices.
 Therefore, the larger $L$ is, the higher the amount of pairs that can be pruned for some specified degradation in the estimation accuracy of the resulting sample covariance.
-The implication is that $K$ in the incomplete U-Statistic can be designed to grow with $O(L)$ instead of $O(L^{2})$, which provides a lot of flexibility for pruning, and will be used for choosing the number of pairs in the next section.
+The implication is that $K$ in the incomplete U-Statistic can be designed to grow with $O(L)$ instead of $O(L^{2})$, which provides considerable flexibility for pruning, eases the overall computational complexity, and will be used for choosing the number of pairs in the next section.
+Once we have determined the incomplete U-statistic formulation of the cross-covariance matrix, we are now in a position to establish the underlying rule of $\mathrm{f}_{1}(\bm{\cdot})$ and $\mathrm{f}_{2}(\bm{\cdot})$ by looking at sample pairwise differences of the confounder.
+Afterwards, we will be employing the pruned cross-covariance matrix to return to kernel methods in the same way as portrayed in (7).
 IV Conditional dependence via U-Statistics
 The conditional cross-covariance matrix between $\bm{\mathsf{u}}$ and $\bm{\mathsf{v}}$ is defined as:
-$\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}|\mathsf{z}}\triangleq\int_{\mathbb{R}}\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}|\mathsf{z}=z}\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z)=\int_{\mathbb{R}}\mathbf{C}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathsf{z}=z}\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z)$ (14)
-where $\mathbf{C}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathsf{z}=z}=\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}|\mathsf{z}=z}$ (given (12)) is the cross-covariance matrix conditioned to a specific value $z$ of a confounder $\mathsf{z}$, and $F_{\mathsf{z}}(z)$ is its cumulative distribution.
-Note that (14) is the most general definition of conditional cross-covariance that is valid for any statistics, while the Schur complement expression is only justified for Gaussian vectors [9, Ch.
-4].
-With the goal of deriving an estimator for (14), and following a similar rationale to $\mathring{\bm{\mathsf{u}}}$ and $\mathring{\bm{\mathsf{v}}}$, we define the virtual random variable $\mathring{\mathsf{z}}\triangleq\mathsf{z}_{1}-\mathsf{z}_{2}$, where $\mathsf{z}_{1}$ and $\mathsf{z}_{2}$ are mutually independent and distributed as $\mathsf{z}$.
-Given that integrating all values of $\mathsf{z}$ is equivalent to doing so for $\mathring{\mathsf{z}}=0$, i.e.
-$\mathsf{z}_{1}=\mathsf{z}_{2}$, the expectation in (14) can be alternatively expressed as
-$\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}|\mathsf{z}}$ $=\int_{\mathbb{R}^{2}}\mathbf{C}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathring{\mathsf{z}}=0}\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z_{1})\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z_{2})$
-$=\mathbf{C}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathring{\mathsf{z}}=0}\int_{\mathbb{R}^{2}}\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z_{1})\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z_{2})=\mathbf{C}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathring{\mathsf{z}}=0},$ (15)
-where $\int_{\mathbb{R}^{2}}\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z_{1})\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z_{2})=1$, and since $\mathbf{C}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathring{\mathsf{z}}=0}$ does not depend on the specific values of $z_{1}$ and $z_{2}$ but rather on them being equal.
-In consequence, conditioning the covariance matrix with respect to $\mathsf{z}$ is equal to conditioning it with respect to $\mathring{\mathsf{z}}=0$.
-Therefore, the result in (IV) suggests letting the potential confounder data control the pruning of the incomplete U-Statistics in (13).
-But, since $\mathring{\mathsf{z}}=0$ is an event of zero probability for continuous random variables, the data control should be based on merely small values of $|\mathring{z}|$.
-Note that the procedure followed here is the same followed in [12].
-With the intention of choosing the data pairs for the pruning based on $|\mathring{z}|$, we define the samples of $\mathring{z}$ as follows:
-$\mathring{z}_{l,l^{\prime}}\triangleq z(l)-z(l^{\prime}),$ (16)
+$\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}|\mathsf{z}}\triangleq\int_{\mathbb{R}}\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}|\mathsf{z}=z}\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z)=\int_{\mathbb{R}}\mathbf{C}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathsf{z}=z}\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z),$ (10)
+where $\mathbf{C}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathsf{z}=z}=\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}|\mathsf{z}=z}$ (given (8)) is the cross-covariance matrix conditioned to a specific value $z$ of confounder $\mathsf{z}$, and $F_{\mathsf{z}}(z)$ is its cumulative distribution.
+With the goal of deriving an estimator of (10), we define the virtual random variable $\mathring{\mathsf{z}}\triangleq\mathsf{z}_{1}-\mathsf{z}_{2}$, where $\mathsf{z}_{1}$ and $\mathsf{z}_{2}$ are mutually independent and distributed as $\mathsf{z}$.
+Integrating over all values of $\mathsf{z}$ is equivalent to doing so over the events in which $\mathsf{z}_{1}$ and $\mathsf{z}_{2}$ take the same value, i.e.
+$\mathring{\mathsf{z}}=0$.
+Therefore, the expectation in (10) can be alternatively expressed as
+$\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}|\mathsf{z}}$ $=\iint_{\mathbb{R}^{2}}\mathbf{C}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathring{\mathsf{z}}=0}\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z_{1})\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z_{2})$
+$=\mathbf{C}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathring{\mathsf{z}}=0}\iint_{\mathbb{R}^{2}}\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z_{1})\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z_{2})=\mathbf{C}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathring{\mathsf{z}}=0},$ (11)
+where $\iint_{\mathbb{R}^{2}}\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z_{1})\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z_{2})=1$, and $\mathbf{C}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathring{\mathsf{z}}=0}$ does not depend on the specific values of $z_{1}$ and $z_{2}$ but rather on them being equal.
+In consequence, conditioning with respect to $\mathsf{z}$ is equal to conditioning with respect to $\mathring{\mathsf{z}}=0$.
+Therefore, (IV) suggests to let the pruning of the incomplete U-Statistics in (9) be handled by the potential confounder.
+However, since $\mathring{\mathsf{z}}=0$ is an event of zero probability for continuous random variables, data control should be based on merely small values of $\lvert\mathring{z}\rvert$ [13].
+With the intention of choosing data pairs to prune according to $\lvert\mathring{z}\rvert$, we define the samples of $\mathring{\mathsf{z}}$ as follows:
+$\mathring{z}_{l,l^{\prime}}\triangleq z(l)-z(l^{\prime}),\qquad l\neq l^{\prime},$ (12)
 where $z(l)$ and $z(l^{\prime})$ are i.i.d.
-samples drawn from $\mathsf{z}$ with $l\neq l^{\prime}$.
-Then, we let the sorting of $|\mathring{z}_{l,l^{\prime}}|$ (in ascending order) be the one that determines the ordering of the index pairs provided by $\mathrm{f}_{1}(\bm{\cdot})$ and $\mathrm{f}_{2}(\bm{\cdot})$ in (11).
-Moreover, in view of Remark 1, the amount of pairs $K$ is set to grow as $O(L)$, i.e.
-$K_{\alpha}\triangleq\bigl{\lfloor}\tfrac{L\alpha}{2}\bigr{\rfloor},$ (17)
+samples drawn from $\mathsf{z}$.
+Then, we let the sorting of $\lvert\mathring{z}_{l,l^{\prime}}\rvert$ (in ascending order) be the one that determines the ordering of the index pairs provided by $\mathrm{f}_{1}(\bm{\cdot})$ and $\mathrm{f}_{2}(\bm{\cdot})$ in Sec.
+III.
+Moreover, in view of Remark 1, the amount of pairs $K$ is set to grow as $O(L)$ with
+$K=\bigl{\lfloor}\tfrac{L\alpha}{2}\bigr{\rfloor},$ (13)
 being $1\leq\alpha\ll(L-1)$ a tuning hyper-parameter.
-While $\alpha=1$ ensures that only very small values of $|\mathring{z}_{l,l^{\prime}}|$ are considered, in the other extreme of $\alpha=L-1$ the U-Statistic becomes complete and there is no conditioning at all, thus encountering the HSIC as a particular case as in (7).
-A natural trade-off then emerges: while $\alpha$ close to $1$ is desirable for the conditioning idea (IV) to work properly, higher values of $\alpha$ will avoid excessive pruning and provide sufficient statistical accuracy in (13) with $K=K_{\alpha}$.
-Remarkably, the selection of the hyper-parameter $\alpha$ becomes a minor issue provided that $L$ is sufficiently large, as was shown in [12] under the correlation measure framework between a pair of vectors, and will be seen later on by a numerical example.
+While $\alpha=1$ ensures that only very small values of $\lvert\mathring{z}_{l,l^{\prime}}\rvert$ are considered, the U-Statistic becomes complete and there is no conditioning at all for $\alpha=L-1$, thus yielding the HSIC as a particular case as in (7).
+Then a natural trade-off emerges: low values of $\alpha$ are desirable to provide strong conditioning of data, but may also lead to excessive pruning with low statistical accuracy in (9).
+Remarkably, the selection of $\alpha$ becomes a minor issue provided that $L$ is sufficiently large, as shown in [13] under the correlation measure framework between a pair of vectors.
+This will be further discussed in Sec.
+V with a numerical example.
 IV-A Conditional HSIC
-Now that we have determined the sorting and pruning of the data pairs according to the confounder $\mathsf{z}$, let us write a conditional dependence measure as the Frobenius norm of (13):
-$\operatorname{\mathrm{tr}}\bigl{(}\widehat{\mathbf{C}}^{\mathrm{H}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathsf{z}}\widehat{\mathbf{C}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathsf{z}}\bigr{)}=\tfrac{1}{K_{\alpha}^{2}}\operatorname{\mathrm{tr}}\bigl{(}\mathring{\mathbf{U}}^{\mathrm{H}}\mathring{\mathbf{U}}\mathring{\mathbf{V}}^{\mathrm{H}}\mathring{\mathbf{V}}\bigr{)},$ (18)
-where the circularity of the trace has been used.
-To link the previous expression with kernel-based methods, let us rewrite the zero-mean virtual data matrices $\mathring{\mathbf{U}}$ and $\mathring{\mathbf{V}}$ as follows:
-$\mathring{\mathbf{U}}=\tfrac{1}{\sqrt{2}}(\mathbf{U}_{1}-\mathbf{U}_{2}),\qquad\mathring{\mathbf{V}}=\tfrac{1}{\sqrt{2}}(\mathbf{V}_{1}-\mathbf{V}_{2}),$ (19)
-where
-$\mathbf{U}_{a}\triangleq[\mathbf{u}_{\mathrm{f}_{a}(1)},\dots,\mathbf{u}_{\mathrm{f}_{a}(K_{\alpha})}],\;\;\mathbf{V}_{a}\triangleq[\mathbf{v}_{\mathrm{f}_{a}(1)},\dots,\mathbf{v}_{\mathrm{f}_{a}(K_{\alpha})}],$ (20)
-for $a=\{1,2\}$. Accordingly, (18) is then rewritten as
-$\operatorname{\mathrm{tr}}\bigl{(}\widehat{\mathbf{C}}^{\mathrm{H}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathsf{z}}\widehat{\mathbf{C}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathsf{z}}\bigr{)}$ (21)
-$=\tfrac{1}{4K_{\alpha}^{2}}\operatorname{\mathrm{tr}}\bigl{(}(\mathbf{U}_{1}-\mathbf{U}_{2})^{\mathrm{H}}(\mathbf{U}_{1}-\mathbf{U}_{2})(\mathbf{V}_{1}-\mathbf{V}_{2})^{\mathrm{H}}(\mathbf{V}_{1}-\mathbf{V}_{2})\bigr{)}.$
-Taking the limit $M\rightarrow\infty$, kernel matrices are obtained from the products among $\mathbf{U}$ and $\mathbf{V}$ in (21):
-$\mathbf{K}_{a,a^{\prime}}\triangleq\lim_{M\rightarrow\infty}\mathbf{U}_{a}^{\mathrm{H}}\mathbf{U}_{a^{\prime}}$ $=\mathbf{S}\mathopen{}\mathclose{{}\left(\mathbf{x}_{a},\mathbf{x}_{a^{\prime}}}\right)\quad$ (22a)
-$\mathbf{Q}_{a,a^{\prime}}\triangleq\lim_{M\rightarrow\infty}\mathbf{V}_{a}^{\mathrm{H}}\mathbf{V}_{a^{\prime}}$ $=\mathbf{S}\mathopen{}\mathclose{{}\left(\mathbf{y}_{a},\mathbf{y}_{a^{\prime}}}\right),$ (22b)
-where $\mathbf{S}\mathopen{}\mathclose{{}\left(\bm{\cdot},\bm{\cdot}}\right)$ is given in (9). Therefore, the new relevant data subsets for inferring conditional dependence are given by
-$[\mathbf{x}_{a}]_{k}\triangleq x(f_{a}(k)),\quad[\mathbf{y}_{a}]_{k}\triangleq y(f_{a}(k)).$ (23)
-Finally, the resulting C-HSIC can be expressed as follows:
+Now that we have determined the sorting and pruning of data pairs according to the confounder $\mathsf{z}$, let us write a conditional dependence measure as the Frobenius norm of (9):
+$\operatorname{\mathrm{tr}}\bigl{(}\widehat{\mathbf{C}}^{\mathrm{H}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathsf{z}}\widehat{\mathbf{C}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathsf{z}}\bigr{)}=\tfrac{1}{K^{2}}\operatorname{\mathrm{tr}}\bigl{(}\mathring{\mathbf{U}}^{\mathrm{H}}\mathring{\mathbf{U}}\mathring{\mathbf{V}}^{\mathrm{H}}\mathring{\mathbf{V}}\bigr{)},$ (14)
+where the circularity of the trace has been used. To link the previous expression with kernel-based methods, let us rewrite the zero-mean virtual data matrix $\mathring{\mathbf{U}}$ as follows:
+$\mathring{\mathbf{U}}=\tfrac{1}{\sqrt{2}}(\mathbf{U}_{1}-\mathbf{U}_{2}),\quad\mathbf{U}_{a}\triangleq[\mathbf{u}_{\mathrm{f}_{a}(1)},\dots,\mathbf{u}_{\mathrm{f}_{a}(K)}],$ (15)
+for $a=\{1,2\}$. The same is done for $\mathring{\mathbf{V}}$. Accordingly, (14) is then rewritten as
+$\operatorname{\mathrm{tr}}\bigl{(}\widehat{\mathbf{C}}^{\mathrm{H}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathsf{z}}\widehat{\mathbf{C}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathsf{z}}\bigr{)}=$ (16)
+$\quad\tfrac{1}{4K^{2}}\operatorname{\mathrm{tr}}\bigl{(}(\mathbf{U}_{1}-\mathbf{U}_{2})^{\mathrm{H}}(\mathbf{U}_{1}-\mathbf{U}_{2})(\mathbf{V}_{1}-\mathbf{V}_{2})^{\mathrm{H}}(\mathbf{V}_{1}-\mathbf{V}_{2})\bigr{)}.$
+Taking the limit of $M\rightarrow\infty$, kernel matrices are obtained from the products among $\mathbf{U}$ and $\mathbf{V}$ in (16):
+$\mathbf{K}_{a,a^{\prime}}\triangleq\lim_{M\rightarrow\infty}\mathbf{U}_{a}^{\mathrm{H}}\mathbf{U}_{a^{\prime}},\qquad\mathbf{Q}_{a,a^{\prime}}\triangleq\lim_{M\rightarrow\infty}\mathbf{V}_{a}^{\mathrm{H}}\mathbf{V}_{a^{\prime}},$ (17)
+whose elements are
+$[\mathbf{K}_{a,a^{\prime}}]_{k,k^{\prime}}$ $=\kappa\bigl{(}x(\mathrm{f}_{a}(k))-x(\mathrm{f}_{a^{\prime}}(k^{\prime}))\bigr{)},$ (18a)
+$[\mathbf{Q}_{a,a^{\prime}}]_{k,k^{\prime}}$ $=\kappa\bigl{(}y(\mathrm{f}_{a}(k))-y(\mathrm{f}_{a^{\prime}}(k^{\prime}))\bigr{)},$ (18b)
+where $\kappa(\bm{\cdot})$ is the kernel function as in (6). Finally, the resulting C-HSIC can be expressed as follows:
 […]
-with $\breve{\mathbf{K}}\triangleq\mathbf{K}_{1,1}+\mathbf{K}_{2,2}-\mathbf{K}_{1,2}-\mathbf{K}_{2,1}$ and $\breve{\mathbf{Q}}\triangleq\mathbf{Q}_{1,1}+\mathbf{Q}_{2,2}-\mathbf{Q}_{1,2}-\mathbf{Q}_{2,1}$.
+with $\breve{\mathbf{K}}\triangleq\mathbf{K}_{1,1}+\mathbf{K}_{2,2}-\mathbf{K}_{1,2}-\mathbf{K}_{1,2}^{\mathrm{H}}$ and $\breve{\mathbf{Q}}\triangleq\mathbf{Q}_{1,1}+\mathbf{Q}_{2,2}-\mathbf{Q}_{1,2}-\mathbf{Q}_{1,2}^{\mathrm{H}}$.
 Note that, as a result of the U-Statistics implementation, each entry of the new kernel-based matrices involves four data samples of the same source ($4$-tuples), in contrast to only the pairs typically involved in classical kernel methods.
@@
 HSIC (7).
-V Numerical illustration
-In order to test the proposed method, we aim at generating uncorrelated data with a controlled amount of co-information [17] (also called interaction information [18]), defined as
-$\mathrm{I}(\mathsf{x};\mathsf{y};\mathsf{z})=\mathrm{I}(\mathsf{x};\mathsf{y})-\mathrm{I}(\mathsf{x};\mathsf{y}|\mathsf{z}),$ (25)
-such that the potentials of the method can be clearly highlighted.
-Similar ideas on modeling conditional dependences can be found in [19].
-Since co-information can be either positive or negative, two scenarios are studied: $\mathcal{M}^{+}$ for $\mathrm{I}(\mathsf{x};\mathsf{y};\mathsf{z})>0$ and $\mathcal{M}^{-}$ for $\mathrm{I}(\mathsf{x};\mathsf{y};\mathsf{z})<0$, whose variables are as follows:
+In contrast to other conditional dependence measures, (19) does not require any matrix inversion.
+Since $K$ is set to grow linearly with $L$ in (13), the computational complexity is $O(L^{2})$.
+V Numerical illustrations
+To test the proposed method, we aim at generating uncorrelated data with a controlled amount of conditional dependence.
+Two scenarios are studied, $\mathcal{M}^{+}$ and $\mathcal{M}^{-}$, defined as follows:
 +:$\mathcal{M}^{+}:\begin{cases}\mathsf{x}=\sqrt{\gamma}\mathsf{a}\mathsf{p}+\mathsf{v}\\ \mathsf{y}=\sqrt{\gamma}\mathsf{a}\mathsf{q}+\mathsf{w}\\ \mathsf{z}=\mathsf{a}\end{cases},\hskip 5.0pt\mathcal{M}^{-}:\begin{cases}\mathsf{x}=\sqrt{\gamma}\mathsf{b}\mathsf{p}+\mathsf{v}\\ \mathsf{y}=\sqrt{\gamma}\mathsf{c}\mathsf{q}+\mathsf{w}\\ \mathsf{z}=\mathsf{b}-\mathsf{c}\end{cases}.$ (20)
@@
 random variables are distributed as $\mathsf{a},\mathsf{b},\mathsf{c}\sim\mathcal{U}(0,\sqrt{3\text{}})$ (uniform), $\mathsf{v},\mathsf{w}\sim\mathcal{N}(0,1)$ (normal), and $\mathsf{p},\mathsf{q}\sim\mathrm{Bern}_{1/2}\{-1,1\}$ (equiprobable Bernoulli).
-Parameter $\gamma$ is the signal-to-noise ratio associated to the measurements and controls the total amount of absolute co-information.
-In model $\mathcal{M}^{+}$, $\mathsf{x}$ and $\mathsf{y}$ are dependent due to the influence of $\mathsf{a}$ in both, but they are conditionally independent, since knowing $\mathsf{z}$ (i.e.
-$\mathsf{a}$) implies that $\mathsf{x}$ and $\mathsf{y}$ become solely driven by independent phenomena ($\mathsf{v}$ and $\mathsf{w}$).
-By contrast, $\mathsf{x}$ and $\mathsf{y}$ are marginally independent in model $\mathcal{M}^{-}$, but they become conditionally dependent, since the knowledge of $\mathsf{z}$ correlates the possible joint values of $\mathsf{b}$ and $\mathsf{c}$.
+In $\mathcal{M}^{+}$, the pair $\{\mathsf{x},\mathsf{y}\}$ are dependent, i.e.
+their mutual information is greater than zero, but they are conditionally independent, since knowing $\mathsf{z}$ implies that $\mathsf{x}$ and $\mathsf{y}$ become solely driven by independent phenomena ($\mathsf{v}$ and $\mathsf{w}$).
+By contrast, $\mathsf{x}$ and $\mathsf{y}$ are marginally independent in $\mathcal{M}^{-}$, but become conditionally dependent, since knowledge of $\mathsf{z}$ correlates the possible joint values of $\mathsf{b}$ and $\mathsf{c}$.
+Parameter $\gamma$ is the signal-to-noise ratio associated with the measurements and controls the total amount of unconditional dependence in $\mathcal{M}^{+}$ and conditional dependence in $\mathcal{M}^{-}$.
 In both models, $\mathsf{x}$, $\mathsf{y}$ and $\mathsf{z}$ are mutually uncorrelated due to the multiplicative effect of mutually independent variables $\mathsf{p}$ and $\mathsf{q}$, so correlation measures are unable to discover any data association.
-Regarding the choice of kernel, the universal Gaussian kernel is used, which yields $\kappa(s)=\exp(-(\frac{s}{\hat{\sigma}L^{-1/5}})^{2})$.
-Note that the kernel length scales in proportion with the sample standard deviation $\hat{\sigma}$ of the data and in inverse proportion with the data size $L$.
-The reader is referred to [7, Appendix D][20, 21] for a justification of the typical power law $O(L^{-1/5})$.
-Figure 1: Model $\mathcal{M}^{+}$.
-Orange: $L=100$; blue: $L=600$.
-Markers denote the empirical average while bands indicate the standard deviation.
-(a)-cross marker: HSIC; (a)-circle marker: C-HSIC; (b): C-HSIC with random pruning.
-Figure 2: Model $\mathcal{M}^{-}$.
-Orange: $L=100$; blue: $L=600$.
-Markers denote the empirical average while bands indicate the standard deviation.
-(a)-cross marker: HSIC; (a)-circle marker: C-HSIC; (b): C-HSIC with random pruning.
-In Fig.
-1 and Fig.
-2, the measures of dependence are shown as a function of $\gamma$ for both models with a moderate hyper-parameter $\alpha=4$ from (17).
-This yields $8\%$ of the total data pairs for $L=100$ and $1.3\%$ for $L=600$ (and even less for higher $L$).
-For model $\mathcal{M}^{+}$ (Fig.
-1), the C-HSIC correctly depicts small conditional dependence, while the HSIC confirms that marginal dependence is high for moderate $\gamma$ values.
+Similar ideas on modeling conditional dependencies can be found in [21], which are inspired by co-information [22, 23].
+The universal Gaussian kernel is used as the kernel function of choice, which yields $\kappa(s)=\exp(-(\frac{s}{\hat{\sigma}L^{-1/5}})^{2})$ being $\hat{\sigma}$ the sample standard deviation.
+We use this expression because of its association with kernel density estimation, known to be related to kernel methods when estimating certain dependence measures [24, Ch.
+2].
+This connection also justifies the adoption of the power law $O(L^{-1/5})$ [25, Ch.
+3][15, Appx.
+D].
+(a) Model $\mathcal{M}^{+}$.
+(b) Model $\mathcal{M}^{-}$.
+Figure 1: Dependence measures as a function of $\gamma$ (with $L=100$) and $L$ (with $\gamma=10$dB).
+C-HSIC measures conditional dependence while HSIC measures unconditional dependence.
+Markers denote the empirical average and bands indicate the standard deviation.
+Fig.
+1 shows measures of dependence for both models with two $\alpha$ values from (13): a moderate $\alpha=4$ and a sixteenfold increase $\alpha=64$.
+This yields $4\%$ and $64.5\%$, respectively, of the total data pairs for $L=100$, or $0.7\%$ and $10.7\%$ for $L=600$ (and even less for higher $L$).
+For $\mathcal{M}^{+}$ (Fig.
+1-(a)), C-HSIC correctly depicts conditional independence for an increasing value of $\gamma$, while HSIC confirms that marginal dependence is high as $\gamma$ increases.
 Conversely, Fig.
-2 exhibits the capability of the C-HSIC to discover significant conditional dependence for moderate $\gamma$ values for model $\mathcal{M}^{-}$, while marginal dependence is confirmed to be small by the HSIC at any value of $\gamma$.
+1-(b) exhibits the capability of C-HSIC to discover conditional dependence for moderate and high $\gamma$ values in $\mathcal{M}^{-}$, while marginal dependence is confirmed to be small by HSIC at any value of $\gamma$.
 To get further insights on the proposed ideas, Fig.
-1-(b) and Fig.
-2-(b) show the performance of the C-HSIC when the pruning of data pairs in the U-Statistics is random, which means that the pair $(l(k),l^{\prime}(k))$ is not controlled by the confounder.
-It is seen that the C-HSIC produces a measure that mimics the marginal dependence one, thus confirming the main property of unbiasedness related to the general U-Statistics theory, and also showing its consequent increase in variance.
-It is also important to note that the selected value of $\alpha$, and the small percentage of data that yields this choice, confirms that the increase in variance induced by the pruning becomes small for sufficiently large data sizes due to Remark 1, making $\alpha$ a noncritical free hyper-parameter.
+1-(a.1) and Fig.
+1-(b.1) also show the performance of C-HSIC when pruning data pairs is performed randomly in the U-statistic, i.e.
+the pair given by $\mathrm{f}_{1}(\bm{\cdot})$ and $\mathrm{f}_{2}(\bm{\cdot})$ is not controlled by the confounder.
+As a result of not being properly conditioned, C-HSIC produces a measure of marginal dependence similar to HSIC, but with increased variance due to the pruning itself.
+It is also worth noting that random and orderly pruning perform similarly for low values of $\gamma$, corresponding to the regime where the confounder has no influence on the pair $\{\mathsf{x},\mathsf{y}\}$, and start to deviate as $\gamma$ increases.
+Additionally, it can be seen that the expected value deteriorates for $\alpha=64$, increasing in $\mathcal{M}^{+}$ and decreasing in $\mathcal{M}^{-}$.
+This is a consequence of a mild pruning for $L=100$ (where $\alpha=64$ is intentionally chosen to illustrate this effect), worsening the conditioning process and starting to behave as an unconditioned measure.
+Finally, Fig.
+1-(a.2) and Fig.
+1-(b.2) illustrate the behavior of C-HSIC for multiple values of $L$ at $\gamma=10$dB.
+Model $\mathcal{M}^{+}$, which is conditionally independent, shows that C-HSIC correctly approaches zero as $L$ increases.
+Conversely, it tends to some nonzero value in $\mathcal{M}^{-}$.
+As mentioned above, $\alpha=64$ is using too many data pairs for small $L$ values and provides a degraded empirical average.
+In both models, $\alpha=4$ shows an increased variance with respect to $\alpha=64$ due to a more thorough pruning.
+However, this is reduced for sufficiently large data sizes due to Remark 1.
+Similarly, the gap between expected values for $\alpha=4$ and $\alpha=64$ is reduced as $L$ increases, albeit the former has the advantage of a lower computational complexity.
+These results show that $\alpha$ is a noncritical hyper-parameter, provided it is reasonably small.
 VI Conclusions
-The classical HSIC measure for marginal statistical dependence has been reinterpreted as a correlation measure on a finite but high-dimensional space based on windowed steering vectors.
-From the original mapping, we can establish an insightful connection with kernels by letting the dimension grow to infinity.
-Instead of being merely a reinterpretation of the HSIC, this vision demonstrates its main power under the more involved conditional dependence framework.
-In particular, the illustrated identification of the HSIC as a sample covariance estimator opens the possibility of leveraging U-Statistics for this task.
-Thanks to this formulation, we can provide a novel route for performing statistical conditioning by pruning data pairs based on the pairwise differences of the confounder under suspicion.
+This work has proposed a proof of concept for a modification on the classical HSIC under the conditional dependence framework, named C-HSIC.
+Its interpretation as a correlation measure on a finite but high-dimensional space allows an insightful connection with kernels by letting the dimension grow without bound.
+Moreover, it opens the possibility of leveraging U-Statistics for this task.
+Thanks to this formulation, we can provide a novel approach for performing statistical conditioning by pruning data pairs based on the pairwise differences of a potential confounder.
 Furthermore, the proposed measure of conditional dependence does not require matrix inversions, which has the advantage of reduced computational complexity and the avoidance of addressing ill-conditioned matrices.
-Further work should study the potential of the proposed method with richer data sets.
+Numerical illustrations have shown that C-HSIC is capable of measuring both conditional dependence and independence in two different scenarios.
+Further research should study the potential of the proposed method with richer data sets, as well as provide a thorough comparison with other methods for measuring conditional dependence.